# ioPUF+: A PUF Based on I/O Pull-Up/Down Resistors for Secret Key Generation in IoT Nodes

### v1

@@ IV Evaluation of PUF Metrics
The first part of this section describes the experimental setup, including the implementation details of ioPUF+ on the PSoC-5 microcontroller. The second part outlines each PUF metric and their evaluation methodology. Finally, we present and discuss the experimental results.
@@ IV-A Experimental Setup
Figure 3 a) shows the schematic of the experimental setup. The setup consists of a PSoC-5 board, CY8CKIT-059 [29], connected to a PC via a USB port. The CY8CKIT-059 board includes two ICs: the PSoC-5, which serves as the device under test, and a UART-over-USB interface used to transfer data from the PSoC-5 to the PC. We used this experimental setup to implement and evaluate four ioPUF+ configurations: (1) ioPUF+ using only the pull-up resistors, (2) ioPUF+ using only the pull-down resistors, (3) ioPUF+ using both pull-up and pull-down resistors, and (4) ioPUF+ using both pull-up and pull-down resistors with SHA-256 hashing.
@@
For configuration (1), we implemented the pull-up resistor measurement circuit shown in Figure 1 b) in the PSoC-5. The implementation used ten I/O pins in resistive pull-up mode ( of ), an analog multiplexer (AMUX), an ADC, and an external resistor, , of . We used the breadboard shown in the figure 3 a) to connect the external pull-down resistor and to wire the output of the MUX to the input of the ADC. We connected the I/O pins sequentially to by configuring the selection bit of the multiplexer in the PSoC-5 firmware. Connecting I/O to resulted in a resistive-divider network, yielding a voltage across , given by:
@@
We used the ADC to measure the voltage . For the measurements, the input range of the ADC was set to 0–5V and the resolution was set to 16 bits, yielding a voltage resolution of approximately . Since we used ten I/O pins, the arrangement facilitated the measurement of ten different voltages, 1 to 10, which are representative of the values of the pull-up resistors corresponding to the ten IOs. We programmed the PSoC-5 to continuously measure 1 to 10 and stream the data via UART at a baud rate of 115200bps to the PC through the UART-Over-USB IC. On the PC side, we used the Tera Term serial terminal to receive and record the data via the PC’s COM port. The PUF responses for the PSoC-5 device was generated from these voltage values on the PC side.
@@
Configuration (2) is similar to configuration (1), with the following differences. For configuration (2), we implemented the pull-down resistor measurement circuit shown in Figure 1 c) in the PSoC-5. We used the breadboard shown in the figure to connect the external pull-up resistor and to wire the output of the MUX to the input of the ADC. In this configuration, voltage values 1 to 10 represented the values of the pull-down resistors corresponding to the ten I/O pins of the PSoC-5. These voltage readings were streamed to the PC. The PUF responses for the PSoC-5 device was generated from these voltage values on the PC side.
@@
Configuration (3) ,as shown in Figure 3 is implemented by combining methods (1) and (2) in the PSoC-5. In this configuration, first, the pull-up resistors of the 10 I/O pins of the PSoC-5 device were measured using the circuit configuration shown in Figure 1 b). The pull-up resistor values were recorded as voltage readings 1 to 10. Next, the pull-down resistors of the 10 IOs of the PSoC-5 were measured using the circuit configuration shown in Figure 1 c). The pull-down resistor values were recorded as voltage readings 11 to 20. These voltage readings, 1 to 20, were then transmitted to the PC. The PUF responses of the PSoC-5 device was generated from these 20 voltage readings on the PC. This configuration acheived PUF responses of length bits, higher than the PUF responses length of bits achieved in configurations (1) and (2).
@@
In all configurations, we recorded voltage measurements from the PSoC-5 device over a 30-minute period. Each set of voltage readings was used to generate a PUF response, resulting in a series of PUF responses for the device over the 30 minutes. These measurements were used to quantify the drift in PUF responses for the device over time. Furthermore, we repeated the experiment on 32 different CY8CKIT-059 boards, i.e., 32 different PSoC-5 devices. The experiments were performed at an ambient temperature of . Figure 3 b) shows the experimental setup with the CY8CKIT-059 connected to the PC.
@@ IV-C1 Reliability, Uniqueness, and Randomness
Figure 4 presents the PUF metrics evaluated for the four ioPUF+ configurations described in Section IV-B. Specifically, the figure includes histogram plots of the intra- and inter-Hamming distances, uniformity, and bit aliasing for the four ioPUF configurations. Table I presents the computed metric values for reliability, uniqueness, uniformity, and bit aliasing.
@@
For Configuration (3) (using both pull-up and pull-down resistors), the intra-Hamming distance remains 0, as expected of an ideal PUF. However, the mean inter-Hamming distance, uniformity, and bit-aliasing deviate from their ideal values, measuring 12.03%, 23.6%, and 23.6%, respectively. These metrics are lower compared to Configurations (1) and (2), as presented in Table I. This behavior likely suggests that combining both resistor types introduces systematic biases in the PUF response generation. Despite these deviations, the non-overlapping distributions of inter and intra-Hamming distances confirm that the PUF responses remain different across the devices as shown in figure 4-3(a).
@@ IV-C2 Stability
Figure 5 (a) shows the experimental setup used to evaluate stability across varying temperatures. The PUF was implemented on the PSoC-5 device of a CY8CKIT-059 board. The board was placed on a hot plate to vary the temperature of the PSoC-5. The die temperature was measured using the PSoC5’s built-in on-chip temperature sensor, which was configured in firmware to report the internal die temperature. Firmware was developed on the PSoC-5 to measure voltages 1 to 20, corresponding to the pull-up and pull-down resistor values, at regular intervals and transmit the data along with die temperature of PSoC-5 to the connected PC via UART over USB.
@@
Figure 5 b) presents the stability of the PUF responses over varying temperatures. The vertical bars represent the minimum and maximum stability values observed across multiple measurements at each temperature point. The results indicate that the PUF responses remains relatively stable across a wide temperature range. The maximum mean stability remained below 1.5%, and the worst-case stability observed was 2.63%, corresponding to five bit flips.
@@
Figure 5 d) presents the stability of the PUF responses across the tested voltage range. The vertical bars represent the minimum and maximum stability values observed across multiple measurements at each voltage setting. The worst-case stability was found to be only 2.10%, corresponding to approximately four bit flips, indicating that the PUF responses remains relatively stable across varying supply voltages.
@@ V-A Secret Key Generation
For generating secret key from PUF responses, we followed the method described in Section III-C. First, the PUF responses were processed using a BCH error-correcting code to generate a unique PUF ID for the PSoC-5. Next, the secret key was derived by applying the SHA-256 hash to the PUF ID.
@@ V-A1 BCH ECC for PUF ID Generation
These requirements were satisfied with and , resulting in the BCH configuration of .
@@
BCH implementation was evaluated on PSoC-5, by manually injecting bit errors into the PUF responses. We noted that the BCH decoder successfully identified and corrected these errors using the helper data from the EEPROM. Figure 6 presents the results of this evaluation.
@@ V-B Encrypting Device-to-Device Communication
In this section, we showcase the use of the PUF-derived secret key to secure device-to-device communication. Specifically, we show the encrypted transfer of ECG data stored in the PSoC-5 to a PC. The PSoC-5 and the PC are physically connected and exchange data via UART over USB. The PSoC-5 stores a few samples of ECG data, which are first encrypted using the secret key and then transferred to the PC. The AES-128 encryption algorithm is used for this demonstration [34]. The ECG signal data is first divided into 16-byte blocks and padded using PKCS#7 when the last block is smaller than 16 bytes. Each block is encrypted using AES-128 in ECB mode. The first 128 bits of the 256 bits secret key is used for AES-128 encryption. The encrypted blocks are transmitted sequentially over UART to the PC via USB interface.
@@
Figure 7(a) presents the schematic of the datapath. Figure 7(b) shows the ECG waveform stored on the PSoC-5, transmitted in encrypted form, and subsequently received and decrypted on the PC. The decrypted waveform is identical to the original hardcoded ECG signal, demonstrating the working of the PUF-based secret key generation and AES encryption framework. Also, the encrypted signal appears random and exhibits no visible characteristics of the original ECG waveform, confirming that the encryption successfully hides the underlying ECG data during transmission.
@@ VI Conclusion
In this work, we present ioPUF+, which introduces a novel PUF that generates unique fingerprints for Integrated Circuits (ICs) utilizing pull-up and pull-down resistor measurements on the I/O pins of these ICs. The method is based on the observation that manufacturing process variations cause the resistor values to differ from their nominal design values in a manner that is unique to each IC. The current PUF design is an extension of our earlier work ioPUF [1]; compared to that, the PUF proposed in this work generates responses of length four-fold higher, i.e., from 45 to 190 bits, and improved performance metrics. The PUF was implemented and characterized on 30 counts of the PSoC-5 microcontroller from Infineon. The work reports detailed characterization of PUF responses, including stability analysis of the responses under varying voltage and supply conditions. The results indicate good PUF performance. Also, beyond just introducing the PUF, ioPUF+ also details the method and implementation of converting PUF responses to derive secret keys using Bose–Chaudhuri–Hocquenghem (BCH) error correction and Secure Hash Algorithm (SHA-256). Finally, an application of ioPUF+ secret keys is demonstrated by securing device-to-device data transfers using AES encryption. A detailed analysis of the resource and power consumption of the complete ioPUF+ system is also reported. The full implementation requires only 19.8 KB of Flash, exhibits a latency of 600 ms, and consumes 79 mW of power, demonstrating the suitability of ioPUF+ for resource-constrained IoT nodes.

### v2

@@ IV Evaluation of PUF Metrics
The first part of this section describes the experimental setup, including the implementation details of ioPUF+ on the PSoC-5 microcontroller. The second part outlines each PUF metric and its evaluation methodology. Finally, we present and discuss the experimental results.
@@ IV-A Experimental Setup
Figure 4 (a) shows the schematic of the experimental setup. The setup consists of a PSoC-5 board, CY8CKIT-059 [29], connected to a PC via a USB port. The CY8CKIT-059 board includes two ICs: the PSoC-5, which serves as the device under test, and a UART-over-USB interface used to transfer data from the PSoC-5 to the PC. We used this experimental setup to implement and evaluate four ioPUF+ configurations: (1) ioPUF+ using only the pull-up resistors, (2) ioPUF+ using only the pull-down resistors, (3) ioPUF+ using both pull-up and pull-down resistors, and (4) ioPUF+ using both pull-up and pull-down resistors with SHA-256 hashing.
@@
For configuration (1), we implemented the pull-up resistor measurement circuit shown in Figure 2 (b) in the PSoC-5. The implementation used ten I/O pins in resistive pull-up mode ( of ), an analog multiplexer (AMUX), an ADC, and an external resistor, , of . We used the breadboard shown in Figure 4 (a) to connect the external pull-down resistor and to wire the output of the MUX to the input of the ADC. We connected the I/O pins sequentially to by configuring the selection bit of the multiplexer in the PSoC-5 firmware. Connecting I/O to resulted in a resistive-divider network, yielding a voltage across , given by:
@@
We used the ADC to measure the voltage . For the measurements, the input range of the ADC was set to 0–5V, and the resolution was set to 16 bits, yielding a voltage resolution of approximately . Since we used ten I/O pins, the arrangement facilitated the measurement of ten different voltages, 1 to 10, which are representative of the values of the pull-up resistors corresponding to the ten I/Os. We programmed the PSoC-5 to continuously measure 1 to 10 and stream the data via UART at a baud rate of 115200bps to the PC through the UART-Over-USB IC. On the PC side, we used the Tera Term serial terminal to receive and record the data via the PC’s COM port. The PUF responses for the PSoC-5 device were generated from these voltage values on the PC side.
@@
Configuration (2) is similar to configuration (1), with the following differences. For configuration (2), we implemented the pull-down resistor measurement circuit shown in Figure 2 (c) in the PSoC-5. We used the breadboard shown in the figure to connect the external pull-up resistor and to wire the output of the MUX to the input of the ADC. In this configuration, voltage values 1 to 10 represented the values of the pull-down resistors corresponding to the ten I/O pins of the PSoC-5. These voltage readings were streamed to the PC. The PUF responses for the PSoC-5 device were generated from these voltage values on the PC side.
@@
Configuration (3), as shown in Figure 4, is implemented by combining methods (1) and (2) in the PSoC-5. In this configuration, first, the pull-up resistors of the 10 I/O pins of the PSoC-5 device were measured using the circuit configuration shown in Figure 2 (b). The pull-up resistor values were recorded as voltage readings 1 to 10. Next, the pull-down resistors of the 10 IOs of the PSoC-5 were measured using the circuit configuration shown in Figure 2 (c). The pull-down resistor values were recorded as voltage readings 11 to 20. These voltage readings, 1 to 20, were then transmitted to the PC. The PUF responses of the PSoC-5 device were generated from these 20 voltage readings on the PC. This configuration achieved PUF responses of length bits, higher than the PUF responses’ length of bits achieved in configurations (1) and (2).
@@
In all configurations, we recorded voltage measurements from the PSoC-5 device over a 30-minute period. Each set of voltage readings was used to generate a PUF response, resulting in a series of PUF responses for the device over the 30 minutes. These measurements were used to quantify the drift in PUF responses for the device over time. Furthermore, we repeated the experiment on 32 different CY8CKIT-059 boards, i.e., 32 different PSoC-5 devices. The experiments were performed at an ambient temperature of . Figure 4 (b) shows the experimental setup with the CY8CKIT-059 connected to the PC.
@@ IV-C1 Reliability, Uniqueness, and Randomness
Figure 5 presents the PUF metrics evaluated for the four ioPUF+ configurations described in Section IV-B. Specifically, the figure includes histogram plots of the intra- and inter-Hamming distances, uniformity, and bit aliasing for the four ioPUF configurations. Table I presents the computed metric values for reliability, uniqueness, uniformity, and bit aliasing. In Table I, the average uniformity and bit-aliasing values are identical, as they compute the same overall fraction of ’1’ bits when averaged across all devices and bit positions, however, their per-device and per-bit distributions differ significantly (see Figure 5).
For Configuration (3) (using both pull-up and pull-down resistors), the intra-Hamming distance remains 0, as expected of an ideal PUF. However, the mean inter-Hamming distance, uniformity, and bit-aliasing deviate from their ideal values, measuring 12.03%, 23.6%, and 23.6%, respectively. These metrics are lower compared to Configurations (1) and (2), as presented in Table I. This behavior suggests that combining both resistor types introduces systematic biases in the PUF response generation. Despite these deviations, the non-overlapping distributions of inter and intra-Hamming distances confirm that the PUF responses remain different across the devices, as shown in figure 5-3(a).
@@ IV-C2 Stability
Figure 6 (a) shows the experimental setup used to evaluate stability across varying temperatures. We implemented the PUF on the PSoC-5 device of a CY8CKIT-059 board. The board was placed on a hot plate to vary the temperature of the PSoC-5. The die temperature was measured using the PSoC-5’s built-in on-chip temperature sensor, which was configured in firmware to report the internal die temperature. Firmware was developed on the PSoC-5 to measure voltages 1 to 20, corresponding to the pull-up and pull-down resistor values, at regular intervals and transmit the data along with the die temperature of PSoC-5 to the connected PC via UART over USB.
@@
Figure 6 (b) presents the stability of the PUF responses over varying temperatures. The vertical bars represent the minimum and maximum stability values observed across multiple measurements at each temperature point. The results indicate that the PUF responses remain relatively stable across a wide temperature range. The maximum mean stability remained below 1.5%, and the worst-case stability observed was 2.63%, corresponding to five bit flips.
@@
Figure 6 (d) presents the stability of the PUF responses across the tested voltage range. The vertical bars represent the minimum and maximum stability values observed across multiple measurements at each voltage setting. The worst-case stability was found to be only 2.10%, corresponding to approximately four bit flips, indicating that the PUF responses remain relatively stable across varying supply voltages.
@@ V-A Secret Key Generation
For generating a secret key from PUF responses, we followed the method described in Section III-C. First, the PUF responses were processed using a BCH error-correcting code to generate a unique PUF ID for the PSoC-5. Next, the secret key was derived by applying the SHA-256 hash to the PUF ID.
@@ V-A1 BCH ECC for PUF ID Generation
These requirements were satisfied with and , resulting in the BCH configuration of . We selected these parameters because they provide sufficient error correction (t=5 bits) for our observed bit error rates while maintaining low computational overhead on the resource-constrained PSoC-5.
@@
BCH implementation was evaluated on PSoC-5 by manually injecting bit errors into the PUF responses. We noted that the BCH decoder successfully identified and corrected these errors using the helper data from the EEPROM. Figure 7 presents the results of this evaluation.
@@ V-B Encrypting Device-to-Device Communication
In this section, we showcase the use of the PUF-derived secret key to secure device-to-device communication. Specifically, we show the encrypted transfer of ECG data stored in the PSoC-5 to a PC. The PSoC-5 and the PC are physically connected and exchange data via UART over USB. The PSoC-5 stores a few samples of ECG data, which are first encrypted using the secret key and then transferred to the PC. The AES-128 encryption algorithm is used for this demonstration [34]. The ECG signal data is first divided into 16-byte blocks and padded using PKCS#7 when the last block is smaller than 16 bytes. Each block is encrypted using AES-128 in ECB mode. The first 128 bits of the 256-bit secret key are used for AES-128 encryption. The encrypted blocks are transmitted sequentially over the UART to the PC via the USB interface.
@@
Figure 8(a) presents the schematic of the datapath. Figure 8(b) shows the ECG waveform stored on the PSoC-5, transmitted in encrypted form, and subsequently received and decrypted on the PC. The decrypted waveform is identical to the original hardcoded ECG signal, demonstrating the working of the PUF-based secret key generation and AES encryption framework. Additionally, the encrypted signal appears random and exhibits no visible characteristics of the original ECG waveform, confirming that the encryption successfully hides the underlying ECG data during transmission.
@@ VI Conclusion
In this work, we present ioPUF+, which introduces a novel PUF that generates unique fingerprints for Integrated Circuits (ICs) utilizing pull-up and pull-down resistor measurements on the I/O pins of these ICs. The method is based on the observation that manufacturing process variations cause the resistor values to differ from their nominal design values in a manner that is unique to each IC. The current PUF design is an extension of our earlier work, ioPUF [1]; compared to that, the PUF proposed in this work generates responses of length fourfold higher, i.e., from 45 to 190 bits, and improved performance metrics. We implemented the PUF and characterized it on 32 PSoC-5 microcontrollers from Infineon. The work reports a detailed characterization of PUF responses, including stability analysis of the responses under varying temperature and voltage conditions. The results indicate good PUF performance. Beyond introducing the PUF, ioPUF+ also details the method and implementation of converting PUF responses to derive secret keys using Bose–Chaudhuri–Hocquenghem (BCH) error correction and Secure Hash Algorithm (SHA-256). Finally, an application of ioPUF+ secret keys is demonstrated by securing device-to-device data transfers using AES encryption. A detailed analysis of the resource and power consumption of the complete ioPUF+ system is also reported. The full implementation requires only 19.8 KB of Flash, exhibits a latency of 600 ms, and consumes 79 mW of power, demonstrating the suitability of ioPUF+ for resource-constrained IoT nodes.